OpenStudy (anonymous):

Joey's Pizza Palace sells four sizes of pizza. The mini pizza has a 6-inch diameter, and the medium pizza has a diameter twice that of the mini pizza. How many more square inches of pizza are there in the medium than in the mini pizza? Round your answer to the nearest tenth.

OpenStudy (lgbasallote):

do you know what twice means?

OpenStudy (lgbasallote):

that awlward moment when everyone viewing the question is replying except the asker lol

OpenStudy (anonymous):

do you the know the formula to calculate the area of a circle?

OpenStudy (anonymous):

multiply the mini pizza by 2 find the area of both and subtract them

OpenStudy (anonymous):

Mini: 6 in. dia. Med.: 6*2= dia. simple. 

OpenStudy (lgbasallote):

the mini pizza has diameter 6..so radius is 3 area of circle = \(\pi r^2 = \pi (3)^2 = 9\pi\) the medium pizza has diameter twice the mini pizza so 12 is diameter..that means 6 is the radius area of circle = \(\pi r^2 = \pi (6^2) = 36\pi\) now it's asking the difference between the two so \(36\pi - 9\pi\)